The sum of the digits of a two digit number is 9. The number with the digits interchanged is 18 more than twice the original number. Find the original number?
Answer by TimothyLamb(4379)

x = ones digit
y = tens digit
---
x + y = 9
10x + y = 2(x + 10y) + 18
---
x + y = 9
10x + y = 2x + 20y + 18
---
x + y = 9
8x - 19y = 18
---
put the system of linear equations into standard form
---
x + y = 9
8x - 19y = 18
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 7
y = 2
---
answer:
the original number is = 27
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php